# Computational Complexity of Physical Information Sufficiency

### v3

@@ 5 Encoding-Regime Separation
The hardness results of Section 4 apply to worst-case instances under the succinct encoding. This section states an encoding-regime separation: an explicit-state upper bound versus a succinct-encoding worst-case lower bound, and an intermediate query-access obstruction family. The complexity class is a property of the encoding measured against a fixed decision question (Remark 2.21), not a property of the question itself: the same sufficiency question admits polynomial-time algorithms under one encoding and worst-case exponential cost under another.
Physical reading: the separation is an access-and-budget separation for the same decision semantics. Explicit access assumes full boundary observability; succinct access assumes compressed structure that may require exponential expansion to certify universally; query access assumes bounded measurement interfaces.[C:T5.1;P5.3,2.34;R:E,S+ETH,Qf,AR] (L: CC16,CC50,CT6.)
@@
Theorem 5.1 has Part 1 in [E] (time polynomial in $|S|$) and Part 2 in [S+ETH] (time exponential in $n$). Proposition 5.3 is [Q_fin] (finite-state Opt-oracle core), while Propositions 5.7 and 5.11 are [Q_bool] interface refinements (value-entry and state-batch). Relative to Definition 2.23, Propositions 5.8 and 5.15 are explicit simulation-transfer instances. The [E] vs [S+ETH] split in Theorem 5.1 is a same-question encoding/cost-model separation, not a bidirectional simulation claim between those regimes. These regimes are defined in Section 2.4 and are not directly comparable as functions of one numeric input-length parameter.
@@
Consequently, no deterministic query procedure using fewer than $|S|$ state queries can solve SUFFICIENCY-CHECK on all such instances; the worst-case query complexity is $\Omega(|S|)$. (L: CC50, L7, L13.)
@@
[C:P5.4;R:H=query-lb] Let $n\geq 1$ and let $P$ be a finite family of Boolean-state pairs. Assume $P$ is sound for refuting empty-set sufficiency in the sense that for every $\text{Opt}:\{0,1\}^{n}\to\{0,1\}$, whenever $\emptyset$ is not sufficient, some pair in $P$ witnesses disagreement of Opt. Then
@@
Equivalently, any sound checker must inspect at least the witness budget $W(n)=2^{n-1}$, so checking budget $T(n)$ satisfies $T(n)\geq W(n)$. (L: DQ587, DQ588.)
@@
[C:C5.5;R:H=query-lb] For the fixed sufficiency question, finite query interfaces induce an information barrier:
@@
[C:C5.6;R:H=query-lb] In the Boolean-coordinate state space $S=\{0,1\}^{n}$, Proposition 5.3 yields the familiar $\Omega(2^{n})$ worst-case query lower bound for Opt-oracle access. (L: CC49; core: L6.)
@@
In the mechanized Boolean value-entry query submodel [Q_bool], for any deterministic procedure using fewer than $2^{n}$ value-entry queries $(a,s)\mapsto U(a,s)$, there exist two queried-value-indistinguishable instances with opposite truth values for SUFFICIENCY-CHECK on the $I=\emptyset$ subproblem. Therefore worst-case value-entry query complexity is also $\Omega(2^{n})$. (L: L9, L19.)
@@
The theorem L2 constructs, for any query set of cardinality $2^{n}$, an unqueried hidden state $s_{0}$ and a yes/no instance pair that agree on all queried values but disagree on $\emptyset$-sufficiency. The auxiliary bound L6 ensures that fewer than $2^{n}$ value-entry queries cannot cover all states, so the indistinguishability argument applies in the worst case.  $\blacksquare$
@@
[C:P5.9;R:H=query-lb] In [Q_bool], for any query set with cardinality $2^{n}$, the indistinguishable yes/no pair from Proposition 5.3 forces one decoding error per random seed for any seed-indexed decoder from oracle transcripts. Consequently, finite-support randomization does not remove the obstruction: averaging preserves a constant error floor on the hard pair. (L: L11, L12, L5.)
@@
[C:P5.10;R:H=query-lb] For any finite-support seed weighting $\mu$, the same hard pair satisfies a weighted identity: the weighted sum of yes-error and no-error equals total seed weight. Hence randomization cannot collapse both errors simultaneously. (L: L23, L24.)
@@
In [Q_bool], the same $\Omega(2^{n})$ lower bound holds for a state-batch oracle that returns the full Boolean-action utility tuple at each queried state. (L: L8, L14.)
@@
[C:P5.12;R:H=query-lb] The empty-subproblem indistinguishability lower-bound core extends from Boolean-vector state spaces to any finite state type with at least two states. (L: L7, L13.)
@@
[C:P5.13;R:H=query-lb] For the const/spike adversary family used in the query lower bounds, querying all states distinguishes the pair; thus the lower-bound family is tight up to constant factors under full-state scan. (L: L10.)
@@
[C:P5.14;R:H=query-lb] Let $w(q)$ be per-query cost and $w_{\min}$ a lower bound on all queried costs. Any cardinality lower bound $|Q|\geq L$ lifts to weighted cost:
@@
In [Q_bool], agreement on state-batch oracle views over touched states implies agreement on all value-entry views for the corresponding entry-query set. (L: L20, CC43.)
@@
‘Opt‘-oracle views are strictly coarser than value-entry views: there exist instances with identical ‘Opt‘ views but distinguishable value entries. (L: L3, L4.)
@@ 5.1 Budgeted Physical Crossover
The [E] vs [S] split can be typed against an explicit physical budget. Let $E(n)$ and $S(n)$ denote explicit and succinct representation sizes for the same decision question at scale parameter $n$, and let $B$ be a hard representation budget. These crossover claims are interpreted through the physical-to-core encoding discipline of Section 2.7: physical assumptions are transferred explicitly into core assumptions, and encoded physical counterexamples map to core failures on the encoded slice.[C:P2.34;C2.35;R:AR] (L: CT6,CT8,CT4.)
@@
[C:P5.18;R:H=exp-lb-conditional] If $E(n)>B$ and $S(n)\leq B$ for some $n$, then explicit representation is infeasible at $(B,n)$ while succinct representation remains feasible at $(B,n)$. (L: CC45, PBC1.)
@@
[C:P5.19;R:H=exp-lb-conditional] Assume there is a global succinct-size cap $C$ and explicit size is unbounded:
@@
[C:P5.20;R:H=exp-lb-conditional] Fix any collapse schema “exact certification competence $\Rightarrow$ complexity collapse” for the target exact-certification predicate. Under the corresponding non-collapse assumption, exact-certification competence is impossible even when a budgeted crossover witness exists. (L: CC46.)
@@
[C:P5.21;R:H=exp-lb-conditional] In the certifying-solver model (Definitions 3.83.9), assume:
@@
[C:C5.22;R:H=exp-lb-conditional] Fix a physical computer model with finite total budget and strictly positive per-bit physical cost. If required operation count for exact certification has an exponential lower bound, then there exists a threshold $n_{0}$ such that for all $n\geq n_{0}$, required physical work exceeds budget. For the canonical requirement model used in this paper, the same eventual-budget-failure conclusion holds. (L: DQ569, DQ568, DQ574.)
@@
[C:P5.23;R:H=exp-lb-conditional] Fix a budgeted encoding model and budget $B$. If a crossover witness exists at budget $B$, then there is a least index $n_{\mathrm{crit}}$ such that crossover holds at $n_{\mathrm{crit}}$, and no smaller index has crossover. (L: DQ576.)
@@
[C:P5.24;R:H=exp-lb-conditional] If explicit size is monotone in $n$ and already exceeds budget $B$ at some $n_{0}$, then explicit infeasibility holds for all $n\geq n_{0}$. (L: DQ577.)
@@
[C:P5.25;R:H=exp-lb-conditional] If, beyond $n_{0}$, explicit encoding is infeasible while succinct encoding remains feasible, then every $n\geq n_{0}$ is a crossover point. (L: DQ575, DQ579.)
@@
[C:C5.26;R:H=exp-lb-conditional] If both explicit and succinct encodings are unbounded, then for every fixed budget $B$, there exists an index where explicit is infeasible and an index where succinct is infeasible. (L: DQ578.)
@@
[C:P5.27;R:H=exp-lb-conditional] Under the same integrity-resource collapse assumption as Proposition 5.21, policy closure holds at any fixed divergence point and uniformly beyond any threshold where crossover holds. (L: DQ580, DQ581.)
5.2 Thermodynamic Lift in Declared Model
The crossover and hardness statements are complexity-theoretic. The thermodynamic lift follows the same assumption discipline as the rest of the paper: claims are exactly those derivable from the declared conversion model and declared bit-operation lower-bound premises.
@@ 5.2 Thermodynamic Lift in Declared Model
[C:P5.28;R:H=cost-growth] Fix a declared thermodynamic conversion model with constants $(\lambda,\rho)$ where $\lambda$ lower-bounds energy per irreversible bit operation and $\rho$ lower-bounds carbon per joule. If a bit-operation lower bound $b_{\mathrm{LB}}\leq b_{\mathrm{used}}$ holds, then:
@@
[C:P5.29;R:H=cost-growth] Under the same non-collapse/collapse schema used by Proposition 3.61, exact-certification competence is impossible; combined with a declared bit-operation lower bound, this yields simultaneous lower bounds on energy and carbon in the declared conversion model. (L: CC67.)
@@
[C:P5.30;R:H=cost-growth] Within the declared linear thermodynamic model, if
@@
[C:P5.31;R:H=cost-growth] Within the same declared linear model, lower-bound accounting is additive over composed bit-operation lower bounds:

### v4

@@ 9 Encoding-Regime Separation
Formally: this section proves regime-indexed access separation for a fixed sufficiency predicate, then lifts bit-operation bounds to thermodynamic/accounting consequences.
The hardness results of Section 4 apply to worst-case instances under the succinct encoding. This section states an encoding-regime separation: an explicit-state upper bound versus a succinct-encoding worst-case lower bound, and an intermediate query-access obstruction family. The complexity class is a property of the encoding measured against a fixed decision question (Remark 2.21), not a property of the question itself: the same sufficiency question admits polynomial-time algorithms under one encoding and worst-case exponential cost under another. Within the paper’s full regime map, this section is the static access-layer decomposition; theorem-level stochastic and sequential core placements are given in Sections 6 and 7, with strict inter-regime separation in Section 8.
Physical reading: the separation is an access-and-budget separation for the same decision semantics. Explicit access assumes full boundary observability; succinct access assumes compressed structure that may require exponential expansion to certify universally; query access assumes bounded measurement interfaces, and the obstruction persists across value-entry/state-batch interfaces with oracle-lattice transfer/strictness closures.[C:T9.1;P9.3, 9.7, 9.11, 9.15, 9.16, 2.42;R:E, S+ETH, Qf, Qb, AR] (L: CC16, CC50, L9, L8, L21, L20, CT6.)
@@
Theorem 9.1 has Part 1 in [E] (time polynomial in $|S|$) and Part 2 in [S+ETH] (time exponential in $n$). Proposition 9.3 is [Q_fin] (finite-state Opt-oracle core), while Propositions 9.7 and 9.11 are [Q_bool] interface refinements (value-entry and state-batch). Relative to Definition 2.29, Propositions 9.8 and 9.15 are explicit simulation-transfer instances. The [E] vs [S+ETH] split in Theorem 9.1 is a same-question encoding/cost-model separation, not a bidirectional simulation claim between those regimes. These regimes are defined in Section 2.4 and are not directly comparable as functions of one numeric input-length parameter.
@@
Consequently, no deterministic query procedure using fewer than $|S|$ state queries can solve SUFFICIENCY-CHECK on all such instances; the worst-case query complexity is $\Omega(|S|)$. (L: CC50, CC50.)
@@
[C:P9.4;R:H=query-lb] Let $n\geq 1$ and let $P$ be a finite family of Boolean-state pairs. Assume $P$ is sound for refuting empty-set sufficiency in the sense that for every $\text{Opt}:\{0,1\}^{n}\to\{0,1\}$, whenever $\emptyset$ is not sufficient, some pair in $P$ witnesses disagreement of Opt. Then
@@
Equivalently, any sound checker must inspect at least the witness budget $W(n)=2^{n-1}$, so checking budget $T(n)$ satisfies $T(n)\geq W(n)$. (L: WD3, WD4, WD5.)
@@
[C:C9.5;R:H=query-lb] For the fixed sufficiency question, finite query interfaces induce an information barrier:
@@
[C:C9.6;R:H=query-lb] In the Boolean-coordinate state space $S=\{0,1\}^{n}$, Proposition 9.3 yields the familiar $\Omega(2^{n})$ worst-case query lower bound for Opt-oracle access. (L: CC49, CC50.)
@@
In the mechanized Boolean value-entry query submodel [Q_bool], for any deterministic procedure using fewer than $2^{n}$ value-entry queries $(a,s)\mapsto U(a,s)$, there exist two queried-value-indistinguishable instances with opposite truth values for SUFFICIENCY-CHECK on the $I=\emptyset$ subproblem. Therefore worst-case value-entry query complexity is also $\Omega(2^{n})$. (L: CC29, CC50.)
@@
The theorem CC50 constructs, for any query set of cardinality $2^{n}$, an unqueried hidden state $s_{0}$ and a yes/no instance pair that agree on all queried values but disagree on $\emptyset$-sufficiency. The auxiliary bound CC50 ensures that fewer than $2^{n}$ value-entry queries cannot cover all states, so the indistinguishability argument applies in the worst case.  $\blacksquare$
@@
[C:P9.9;R:H=query-lb] In [Q_bool], for any query set with cardinality $2^{n}$, the indistinguishable yes/no pair from Proposition 9.3 forces one decoding error per random seed for any seed-indexed decoder from oracle transcripts. Consequently, finite-support randomization does not remove the obstruction: averaging preserves a constant error floor on the hard pair. (L: CC50, CC50.)
@@
[C:P9.10;R:H=query-lb] For any finite-support seed weighting $\mu$, the same hard pair satisfies a weighted identity: the weighted sum of yes-error and no-error equals total seed weight. Hence randomization cannot collapse both errors simultaneously. (L: CC50, CC50.)
@@
In [Q_bool], the same $\Omega(2^{n})$ lower bound holds for a state-batch oracle that returns the full Boolean-action utility tuple at each queried state. (L: CC28, CC50.)
@@
[C:P9.12;R:H=query-lb] The empty-subproblem indistinguishability lower-bound core extends from Boolean-vector state spaces to any finite state type with at least two states. (L: CC50, CC50.)
@@
[C:P9.13;R:H=query-lb] For the const/spike adversary family used in the query lower bounds, querying all states distinguishes the pair; thus the lower-bound family is tight up to constant factors under full-state scan. (L: CC50.)
@@
[C:P9.14;R:H=query-lb] Let $w(q)$ be per-query cost and $w_{\min}$ a lower bound on all queried costs. Any cardinality lower bound $|Q|\geq L$ lifts to weighted cost:
@@
In [Q_bool], agreement on state-batch oracle views over touched states implies agreement on all value-entry views for the corresponding entry-query set. (L: CC43, CC52.)
@@
‘Opt‘-oracle views are strictly coarser than value-entry views: there exist instances with identical ‘Opt‘ views but distinguishable value entries. (L: CC27, CC29.)
@@
Informally: if you can see enough of the decision boundary, checking can be fast; with limited access, it becomes expensive.
@@ 9.1 Budgeted Physical Crossover
Formally: this subsection states budget-indexed representational feasibility splits and links them to integrity-preserving policy closure.
The [E] vs [S] split can be typed against an explicit physical budget. Let $E(n)$ and $S(n)$ denote explicit and succinct representation sizes for the same decision question at scale parameter $n$, and let $B$ be a hard representation budget. These crossover claims are interpreted through the physical-to-core encoding discipline of Section 2.7: physical assumptions are transferred explicitly into core assumptions, and encoded physical counterexamples map to core failures on the encoded slice.[C:P2.42;C2.43;R:AR] (L: CT6, CT8, CT4.)
@@
[C:P9.18;R:H=exp-lb-conditional] If $E(n)>B$ and $S(n)\leq B$ for some $n$, then explicit representation is infeasible at $(B,n)$ while succinct representation remains feasible at $(B,n)$. (L: CC45, PBC1.)
@@
[C:P9.19;R:H=exp-lb-conditional] Assume there is a global succinct-size cap $C$ and explicit size is unbounded:
@@
[C:P9.20;R:H=exp-lb-conditional] Fix any collapse schema “exact certification competence $\Rightarrow$ complexity collapse” for the target exact-certification predicate. Under the corresponding non-collapse assumption, exact-certification competence is impossible even when a budgeted crossover witness exists. (L: CC46.)
@@
The same closure form is isolated as a standalone physical-collapse chain: requirement-indexed collapse schemas (PH11) induce no physical collapse at that requirement (PH12); canonical specialization yields P=NPP{=}NP physical impossibility under the declared collapse map (PH13, PH14, PH15). The collapse map itself is derived in an explicit SAT bridge (polytime SAT witness + reduction bridge + runtime-to-energy transfer) (PH16, PH17, PH18, PH19, PH20, PH21, PH22), and the assumption boundary is explicit: under bounded budget, positive per-bit cost, and exponential required-op lower bound, claimed collapse forces at least one assumption failure (PH26, PH27); finite-budget necessity is witnessed by explicit unbounded-budget collapse countermodels (PH31, PH32, PH33).
@@
[C:P9.21;R:H=exp-lb-conditional] In the certifying-solver model (Definitions 3.83.9), assume:
@@
[C:C9.22;R:H=exp-lb-conditional] Fix a physical computer model with finite total budget and strictly positive per-bit physical cost. If required operation count for exact certification has an exponential lower bound, then there exists a threshold $n_{0}$ such that for all $n\geq n_{0}$, required physical work exceeds budget. For the canonical requirement model used in this paper, the same eventual-budget-failure conclusion holds. (L: PBC7, PBC8.)
@@
[C:P9.23;R:H=exp-lb-conditional] Fix a budgeted encoding model and budget $B$. If a crossover witness exists at budget $B$, then there is a least index $n_{\mathrm{crit}}$ such that crossover holds at $n_{\mathrm{crit}}$, and no smaller index has crossover. (L: PBC5.)
@@
[C:P9.24;R:H=exp-lb-conditional] If explicit size is monotone in $n$ and already exceeds budget $B$ at some $n_{0}$, then explicit infeasibility holds for all $n\geq n_{0}$. (L: PBC7.)
@@
[C:P9.25;R:H=exp-lb-conditional] If, beyond $n_{0}$, explicit encoding is infeasible while succinct encoding remains feasible, then every $n\geq n_{0}$ is a crossover point. (L: PBC9, PBC8.)
@@
[C:C9.26;R:H=exp-lb-conditional] If both explicit and succinct encodings are unbounded, then for every fixed budget $B$, there exists an index where explicit is infeasible and an index where succinct is infeasible. (L: PBC10.)
@@
[C:P9.27;R:H=exp-lb-conditional] Under the same integrity-resource collapse assumption as Proposition 9.21, policy closure holds at any fixed divergence point and uniformly beyond any threshold where crossover holds. (L: PBC11, PBC12.)
Informally: representational crossover can happen even when exact certification is blocked; under integrity this means abstention or reduced coverage.
9.2 Thermodynamic Lift and Cost Accounting
Formally: this subsection converts bit-operation lower bounds into energy/carbon lower bounds via linear conversion and additive accounting.
The crossover and hardness statements are complexity-theoretic. The thermodynamic lift follows the same assumption discipline as the rest of the paper: claims are exactly those derivable from the conversion model and bit-operation lower-bound premises.
@@ 9.2 Thermodynamic Lift and Cost Accounting
[C:P9.28;R:H=cost-growth] Fix a thermodynamic conversion model with constants $(\lambda,\rho)$ where $\lambda$ lower-bounds energy per irreversible bit operation and $\rho$ lower-bounds carbon per joule. If a bit-operation lower bound $b_{\mathrm{LB}}\leq b_{\mathrm{used}}$ holds, then:
@@
[C:P9.29;R:H=cost-growth] Under the same non-collapse/collapse schema used by Proposition 3.61, exact-certification competence is impossible; combined with a declared bit-operation lower bound, this yields simultaneous lower bounds on energy and carbon in the declared conversion model. (L: CC67.)
@@
[C:P9.30;R:H=cost-growth] Within the linear thermodynamic model, if
@@
[C:P9.31;R:H=cost-growth] Within the same linear model, lower-bound accounting is additive over composed bit-operation lower bounds: